# Universal neural network potentials as descriptors: Towards scalable chemical property prediction using quantum and classical computers

### v1

@@
Abstract
@@
Accurate prediction of diverse chemical properties is crucial for advancing molecular design and materials discovery. Here we present a versatile approach that uses the intermediate information of a universal neural network potential as a general-purpose descriptor for chemical property prediction. Our method is based on the insight that by training a sophisticated neural network architecture for universal force fields, it learns transferable representations of atomic environments. We show that transfer learning with a graph neural network potential M3GNet achieves accuracy comparable to state-of-the-art methods for predicting the NMR chemical shifts ofH, C, N, O, and F using quantum machine learning as well as a standard classical regression model, despite the compactness of its descriptors. This work provides an efficient way to accurately predict properties, potentially accelerating the discovery of new molecules and materials.
I Introduction
@@
As evidenced by the enumeration of 166.4B possible organic molecules containing up to 17 heavy elements, such as C, N, O, S, and halogens (excluding hydrogen), the expansion of the chemical space is astronomical with the increase in types and numbers of elements.[1, 2] This vast landscape has given rise to a multidisciplinary approaches to combining experimental and computational chemistry for the discovery of new chemical substances and materials in a wide range of fields, including material, catalysis, and drug design.[1, 2, 3, 4, 5, 6] Although quantum chemistry and first-principles calculations offer accurate descriptions of chemical substances, their high computational demands make an exhaustive exploration of the chemical space impractical.[7, 8, 9, 10, 11, 12, 6] However, machine- and deep-learning techniques are overcoming these limitations to enable a more extensive exploration.[13, 14, 15, 16, 17, 10, 18, 19, 20, 21, 4, 22, 23, 24, 25, 26, 27, 6]
@@
With machine learning, physics-inspired descriptors that characterize the chemical space have been developed and serve as the cornerstone for building efficient and highly accurate models.[28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 21, 38, 23, 39, 40] Smooth overlap of atomic positions (SOAP)[28, 14, 41, 31, 32, 18, 21], Faber–Christensen–Huang–Lilienfeld (FCHL)[29, 30, 33, 34, 38], and similar descriptors offer atom-level descriptions within molecular or material environments based on physical insights and are effective in regressing chemical quantities, such as interatomic potentials (IAP) and nuclear magnetic resonance (NMR) chemical shifts.[28, 14, 41, 42, 29, 30, 15, 31, 32, 33, 34, 35, 36, 11, 37, 21, 38, 23, 24, 43] Notably, IAPs built using descriptors and Gaussian Process Regression (GPR) [14] have been termed Gaussian approximation potentials (GAP) and have found success in the exploration of the chemical space of molecules and materials.[14, 37, 21] Both kernel ridge regression (KRR) and GPR have been employed to improve the accuracy of NMR chemical shift prediction [44, 41, 42, 29, 30, 43]. However, the dimensionality of the descriptors becomes a barrier to generalization and high accuracy as the molecular or material composition becomes more diverse owing to the addition of different types of elements.[35, 19, 45, 39]
@@
Recently, deep-learning models based on graph neural networks (GNNs) have been proposed to describe chemical spaces using graph representations.[46, 47, 48, 49, 50, 9, 17, 18, 51, 52, 19, 53, 20, 45, 54, 55, 56, 24, 25, 57] In most GNN-based IAPs, atoms within a molecular or material environment are represented as nodes, and their local connectivity as edges in a graph. The graph is then convolved to embed atom-specific information within each node, and further processed using multilayer perceptrons (MLP) to predict target observables. Numerous high-quality GNN-based IAPs have been developed for screening high-accuracy first-principles calculations and proposing molecular and crystal structures. Universal graph-based IAPs capable of comprehensively handling molecules and materials, such as MEGNet[52], M3GNet[10], Allegro[58], GNoME[59], and MACE[60, 61, 62] have emerged and are advancing rapidly.
@@
Similarly, GNN-based models have been developed to predict NMR chemical shifts.[46, 47, 49, 50, 53, 57] For example, Yanfei Guan et al. achieved an accuracy of for H and for C by training the SchNet architecture [51] on molecular NMR chemical shifts (CASCADE)[46]. However, the scalability remains an issue due to the increasing optimization costs of GNN and MLP parameters when the size of datasets increase. Han et al. addressed this issue by constraining the nodes in a GNN to heavy elements only, thereby rendering the construction of scalable GNN-based NMR chemical shift models feasible while achieving a state-of-the-art prediction accuracy comparable to that of CASCADE.[53]
@@
Notably, both descriptor-based and GNN methods face challenges. The former faces increased learning costs as the composition becomes more complex, and the latter faces increasing parameter optimization costs with larger training datasets. To address these issues simultaneously, we focused on the potential utility of the outputs from pre-trained GNN-based IAPs as descriptors. We considered these outputs GNN transfer learning (GNN-TF) descriptors and built machine-learning models for predicting chemical properties. Note that there are existing studies attempting to apply pre-trained GNN potentials to other tasks, particularly to generative modeling. [63, 64, 65, 66]
@@
The remainder of this paper is organized as follows. Section II details the GNN-TF descriptor and the kernel method, implemented on both classical and quantum computers, for predicting NMR chemical shifts of H, C, N, O, and F. Section III presents the performance of our developed machine learning models. Section IV discusses the benefits and applications of the GNN-TF descriptor. Finally, Section V concludes the paper.
II Method:
@@
Transfer Learning Using Pre-trained Graph Neural Network
In this section, we discuss the transfer learning of a pre-trained GNN-based IAP. This approach integrates the outputs from the GNN layer of the IAP as shown in Fig. 1.
The architecture of a GNN-based IAP can be broadly segmented into a GNN layer and an MLP layer (gray area of Fig. 1). For the IAP, we opted for two backbones: a MEGNet pre-trained on the QM9 dataset[67] and a M3GNet trained on the MPF.2021.2.8 dataset, which encompasses compounds covering all 89 elements from the Materials Project.[10]
The parameters of the GNN layer in the M3GNet IAP have been optimized for the prediction of system energy, forces, and stress tensors. When fed with the coordinates, s, of a molecule with atoms, the GNN layer generates a vector, , which mirrors the environment of the th atom in the molecule—henceforth referred to as the GNN-TF descriptor.
The GNN layer for both MEGNet and M3GNet outputs GNN-TF descriptors with dimensions of 32 and 64 per atom, respectively. Using GNN-TF descriptors as input, a regression model was constructed to predict NMR chemical shielding constants.
@@
For the regressor, one can choose methodologies, such as GPR, KRR, or feed-forward neural network (NNs), which are contingent on the specific task. In this study, to ensure a maximally fair comparison with other descriptor-based techniques, we adopted KRR.
@@
KRR combines the merits of ridge regression, which offers regularization to mitigate overfitting, with the kernel method, facilitating nonlinear regression. In kernel methods, the data —in the context of our study, the GNN-TF descriptors— are mapped into a high-dimensional feature space through a non-linear kernel function. The Laplacian kernel, which is known to produce favorable results for chemical properties prediction [68, 13, 29], was employed
@@
where represents the hyperparameter of the kernel. In KRR, the predicted value for the target chemical property of the target atom is derived from the GNN-TF descriptor as follows:
@@
All computations related to the KRR were executed using Scikit-learn v.1.2.2,[69] and the hyperparameters of each model were tuned using Optuna v.2.10.[70] The hyperparameters of all models were obtained using only one random 80:20 split of the dataset. We employed five-fold cross-validation for all hyperparameter optimizations. For dataset sizes of up to 50K items, we conducted hyperparameter optimization for 100 iterations, whereas for datasets containing 100K items, we limited the optimization to 50 iterations.
@@
The quantum-kernel method leverages quantum computers to compute kernels, [16, 71, 72] which is achieved by embedding feature vectors generated by classical computers into quantum states. This method calculates the inner product of these quantum states to derive the desired kernels. Embedding feature vectors into quantum states corresponds to mapping them onto a Hilbert space with dimensions raised to the power of two quantum bits (qubits). Using the kernel matrix constructed on a quantum computer, we performed a quantum kernel ridge regression (KRR), commonly referred to as quantum KRR (QKRR).
@@
In this study, we adopted the natural parameterized quantum circuits (NPQC) Kernel, which has been demonstrated to possess performance characteristics similar to the Gaussian kernel, both theoretically and in actual hardware experiments [73, 74, 75]. All computations were conducted using Scikit-qulacs [69, 76, 77]. The quantum kernel was constructed in a 10-qubit space. Hyperparameters for the quantum kernel were determined through grid search. The determined parameters of NPQC kernel were = 1.5 and the repetition times of embedding 40. The regularization hyperparameter in QKRR was determined using 10 iterations of randomized search.
III Results
@@
In Section III.1, because we deal with many elements, we compared the dimensional efficiency of our proposed GNN-TF descriptor to well-established physics-inspired descriptors. Note that the GNN-TF descriptor can better handle complex chemical systems by exploiting the GNN-based IAP architecture.
@@
In Section III.2, we focused on the accuracy of the GNN-TF descriptor in predicting NMR chemical shifts, which are key to understanding molecular details (e.g., interatomic distances and bond angles). This scenario provides an ideal test for determining how well the GNN-TF descriptor works in our study.
@@
Our analysis began by comparing quantum kernel learning, in which the kernels are tested using a quantum computer emulator with traditional kernel learning methods. We then checked the accuracy of the GNN-TF descriptors across the different pretrained GNN models.
@@
Finally, we juxtaposed our GNN-TF descriptor using well-established physics-inspired descriptors. This comparison demonstrates the superiority of the proposed descriptor in terms of efficiency and accuracy. Furthermore, it highlights its potential for accurately predicting chemical properties, which is crucial for advancing research in the molecular and material sciences.
III.1 Dimensional Efficiency
| SOAP111SOAP and FCHL were generated by Dscribe 0.4.0[28] and QML 0.4.0.12[78], respectively. The default hyperparameters were selected as in QM9NMR paper. | FCHL1911footnotemark: 1 |
| Scaling with respect to | |||||
| 5740 | 740 | 64 | |||
| 22680 | 2440 | 64 | |||
| 1737120 | 162336 | 64 |
At the atomic level, descriptors are tools designed to encode information about atoms within molecules or crystalline materials into vectors. Popular descriptors, such as SOAP and FCHL18, excel at intricately capturing the environment within an atom’s cutoff radius.
@@ III.1 Dimensional Efficiency
Although these descriptors have achieved significant success in various accuracy benchmarks, they also present challenges due to their large dimensions. Various strategies have been developed to address these challenges,[79, 80, 81, 34] including refining the descriptor itself, using principal component analysis for dimensionality reduction, and exploring NNs to encode them. In particular, Christensen et al. applied Behler’s method of the atom-centered symmetry function [82] for NN potential to discretize FCHL18[33] to derive a compact and accurate FCHL19.[34]
In Table I, we present the scaling of the SOAP, FCHL19, and M3GNet GNN-TF descriptors in response to an increase in the number of elemental species considered. Additionally, for the QM9, QMugs,[83] and MPF.2021.8 datasets,[10] the descriptor dimensions corresponding to 5, 10, and 89 elemental species comprising each dataset are summarized, respectively.
@@
Remarkably, with an increase in the number of element types, both SOAP and FCHL19 exhibited quadratic scaling. As a snapshot, when representing five elements in the QM9 dataset, the SOAP and FCHL19 methods have dimensions of 5,740 and 740, respectively. This dimensional disparity increases with the number of elemental types. Hence, to represent the 89 elements, the dimensions increased to 1,737,120 and 162,336, respectively. This is tens-of-thousands times larger than the concise 64 dimensions of the M3GNet GNN-TF descriptor.
Owing to its consistent dimensionality, irrespective of the increase in elements, the GNN-TF descriptor is overwhelmingly compact.
III.2 Prediction Accuracy: NMR Chemical Shifts

### v2

# ![[Uncaptioned image]](x3.png) Universal neural network potentials as descriptors: Towards scalable chemical property prediction using quantum and classical computers | | Tomoya Shiota,∗a,b Kenji Ishihara,b and Wataru Mizukami∗a,b | |

@@
![[Uncaptioned image]](x3.png)
Universal neural network potentials as descriptors: Towards scalable chemical property prediction using quantum and classical computers |
| Tomoya Shiota,∗a,b Kenji Ishihara,b and Wataru Mizukami∗a,b | |
@@
Accurate prediction of diverse chemical properties is crucial for advancing molecular design and materials discovery. Here we present a versatile approach that uses the intermediate information of a universal neural network potential as a general-purpose descriptor for chemical property prediction. Our method is based on the insight that by training a sophisticated neural network architecture for universal force fields, it learns transferable representations of atomic environments. We show that transfer learning with graph neural network potentials such as M3GNet and MACE achieves accuracy comparable to state-of-the-art methods for predicting the NMR chemical shifts of using quantum machine learning as well as a standard classical regression model, despite the compactness of its descriptors. In particular, the MACE descriptor demonstrates the highest accuracy to date on the 13C NMR chemical shift benchmarks for drug molecules. This work provides an efficient way to accurately predict properties, potentially accelerating the discovery of new molecules and materials. |
1 Introduction
@@
As evidenced by the enumeration of 166.4B possible organic molecules containing up to 17 heavy elements, such as C, N, O, S, and halogens (excluding hydrogen), the expansion of the chemical space is astronomical with the increase in types and numbers of elements.1, 2 This vast landscape has given rise to multidisciplinary approaches to combining experimental and computational chemistry for the discovery of new chemical substances and materials in a wide range of fields, including material, catalysis, and drug design.1, 2, 3, 4, 5, 6 Although quantum chemistry and first-principles calculations offer accurate descriptions of chemical substances, their high computational demands make an exhaustive exploration of the chemical space impractical.7, 8, 9, 10, 11, 12, 6 However, machine- and deep-learning techniques are overcoming these limitations to enable a more extensive exploration.13, 14, 15, 16, 17, 10, 18, 19, 20, 21, 4, 22, 23, 24, 25, 26, 27, 6
@@
With machine learning, physics-inspired descriptors that characterize the chemical space have been developed and serve as the cornerstone for building efficient and highly accurate models.28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 21, 38, 23, 39, 40 Smooth overlap of atomic positions (SOAP)28, 14, 41, 31, 32, 18, 21, Faber–Christensen–Huang–Lilienfeld (FCHL)29, 30, 33, 34, 38, and similar descriptors offer atom-level descriptions within molecular or material environments based on physical insights and are effective in regressing chemical quantities, such as interatomic potentials (IAP) and nuclear magnetic resonance (NMR) chemical shifts.28, 14, 41, 42, 29, 30, 15, 31, 32, 33, 34, 35, 36, 11, 37, 21, 38, 23, 24, 43 Notably, IAPs built using descriptors and Gaussian process regression (GPR) 14 have been termed Gaussian approximation potentials (GAP) and have found success in the exploration of the chemical space of molecules and materials.14, 37, 21 Both kernel ridge regression (KRR) and GPR have been employed to improve the accuracy of NMR chemical shift prediction 44, 41, 42, 29, 30, 43. However, the dimensionality of the descriptors becomes a barrier to generalization and high accuracy as the molecular or material composition becomes more diverse owing to the addition of different types of elements.35, 19, 45, 39
@@
Recently, deep-learning models based on graph neural networks (GNNs) have been proposed to describe chemical spaces using graph representations.46, 47, 48, 49, 50, 9, 17, 18, 51, 52, 19, 53, 54, 55, 20, 45, 56, 57, 58, 24, 25, 59, 60, 10, 61, 62, 63, 64 In most GNN-based IAPs, atoms within a molecular or material environment are represented as nodes, and their local connectivity as edges in a graph. The graph is then convolved to embed atom-specific information within each node, and further processed using multilayer perceptrons (MLP) to predict target observables. In molecular and materials simulation and modeling, the consideration of symmetry is extremely important. It is desirable for GNNs to be invariant or equivariant to symmetry operations such as translation, rotation, and reflection for the models to make physically meaningful predictions. GNNs that possess these properties are referred to as invariant GNNs or equivariant GNNs. The universal GNN-based IAPs proposed thus far have been designed to satisfy these symmetries. Recently, E(3) or SE(3) equivariant GNN-based IAPs (e.g., Allegro61, GNoME65, MACE62, 63, 64) have demonstrated superior performance compared to E(3) invariant GNN-based IAPs (e.g., MEGNet52, M3GNet10).66, 67
Similarly, GNN-based models have been developed to predict NMR chemical shifts.46, 47, 49, 50, 68, 59 DFT-level calculations of NMR chemical shifts for 1H and 13C have demonstrated the ability to predict within a target accuracy range of 1-2% relative to the possible ranges of approximately 10 ppm and 200 ppm, respectively69, 70. Therefore, the uncertainty in machine learning models using DFT-level datasets is this level of precision, with the target accuracy of 0.2 ppm for 1H and 2 ppm for 13C.30 For example, Yanfei Guan et al. achieved the target accuracy of for 1H and for 13C by training the SchNet architecture 51 on molecular NMR chemical shifts (CASCADE)46.
@@
However, the scalability remains an issue due to the increasing optimization costs of GNN and MLP parameters when the size of datasets increase. Han et al. addressed this issue by constraining the nodes in a GNN to heavy elements only, thereby rendering the construction of scalable GNN-based NMR chemical shift models feasible while achieving a state-of-the-art prediction accuracy comparable to that of CASCADE.68 Furthermore, NMR chemical shifts of various nuclei beyond hydrogen and carbon have become crucial for understanding systems involving a wide range of elements, such as proteins and solids.71, 72, 73, 74, 75, 76, 77 Consequently, efforts are being made to develop machine learning models for NMR chemical shifts of nuclei such as 15N, 17O, and 19F. 75, 76, 73, 74 These elements exhibit wide chemical shift ranges, with about 600, 2500, 500 ppm for 15N, 17O, and 19F, respectively. The target accuracy for these nuclei is set at 25 ppm for 15N and 5 ppm for 19F as well as 1H and 13C.71, 72, 73, 74, 75, 76, 77
@@
Notably, both descriptor-based and GNN-based methods face challenges. The former faces increased learning costs as the composition becomes more complex, and the latter faces increasing parameter optimization costs with larger training datasets. To address these issues simultaneously, we focused on the potential utility of the outputs from pre-trained GNN-based IAPs as descriptors. We considered these outputs GNN transfer learning (GNN-TL) descriptors and built machine-learning models for predicting chemical properties. Note that there are existing studies attempting to apply pre-trained GNN potentials to other tasks, particularly to generative modeling. 78, 79, 80, 81
@@
The remainder of this paper is organized as follows. Section 2 details the GNN-TL descriptor and the kernel method, implemented on both classical and quantum computers, for predicting NMR chemical shifts of 1H, 13C, 15N, 17O, and 19F. Section 3 presents the performance of our developed machine learning models. Section 4 discusses the benefits and applications of the GNN-TL descriptor. Finally, Section 5 concludes the paper.
@@
2 Method: Transfer Learning Using Pre-trained Graph Neural Network
In this section, we discuss the transfer learning of a pre-trained GNN-based IAP. This approach integrates the outputs from the GNN layer of the IAP as shown in Fig. 1. The architecture of a GNN-based IAP can be broadly segmented into a GNN layer and an MLP layer (gray area of Fig. 1). For the E(3) invariant GNN-based IAP, we opted for two backbones: a MEGNet pre-trained on the QM9 dataset82 and a M3GNet trained on the MPF.2021.2.8 dataset, which encompasses compounds covering all 89 elements from the Materials Project.10 The parameters of the GNN layer in the M3GNet IAP were optimized to predict system energy, forces, and stress tensors. Additionally, we incorporated the E(3) equivariant GNN-based IAPs, namely MACE62, 64, 63, into our study. We employed two types of pre-trained MACE IAPs: one trained on a larger dataset named MPtrj83 from Materials Project, referred to as the MACE-MP0 model64, and another trained on an organic molecule dataset covering 10 types of elements including SPICE84 and QMug85, termed the MACE-OFF23 model63. Each model has variations in parameter size, and for this study, we utilized the "small" and "large" versions.64, 63
When fed with the atomic coordinates of a molecule with atoms, denoted by , where represents the atomic number indicating the type of each atom, and is the three-dimensional position vector of the th atom, the GNN layer generates a set of vectors, , which mirrors the environment of the th atom in the molecule. This is referred to as the GNN-TL descriptor. The GNN layer for both MEGNet and M3GNet outputs GNN-TL descriptors with dimensions of 32 and 64 per atom, respectively. On the other hand, MACE is a GNN architecture that predicts energy in the form of atomic cluster expansion. As in Ref.86, only the output of the 1st layer of the GNN layer, corresponding to the one-body term of the many-body expansion, is used as the GNN-TL descriptor. The dimensions of this GNN-TL descriptor are 128, 256, 96, and 224 per atom for MACE-MP0-small, MACE-MP0-large, MACE-OFF23-small, and MACE-OFF23-large, respectively.
@@ 2 Method: Transfer Learning Using Pre-trained Graph Neural Network
Using GNN-TL descriptors as input, a regression model was constructed to predict NMR chemical shielding constants. For the regressor, one can choose methodologies, such as GPR, KRR, or feed-forward neural network (NNs), which are contingent on the specific task. In this study, to ensure a maximally fair comparison with other descriptor-based techniques, we adopted KRR.
@@
KRR combines the merits of ridge regression, which offers regularization to mitigate overfitting, with the kernel method, facilitating nonlinear regression. In kernel methods, the data —in the context of our study, the GNN-TL descriptors— are mapped into a high-dimensional feature space through a non-linear kernel function. The Laplacian and Gaussian kernels were applied:
@@
where is the hyperparameter of the kernel and is the norm parameter that differentiates the type of kernel: for the Laplacian kernel and for the Gaussian kernel. In KRR, the predicted value for the target chemical property of the target atom is derived from the GNN-TL descriptor as follows:
@@
All computations related to the KRR were executed using Scikit-learn v.1.2.2,87 and the hyperparameters of each model were tuned using Optuna v.2.10.88 For dataset sizes of up to 50K items, we conducted hyperparameter optimization for 100 iterations with ten-fold cross-validation, while for those at 100K, we limited the optimization to 10 iterations.
@@
The quantum-kernel method leverages quantum computers to compute kernels, 16, 89, 90 which is achieved by embedding feature vectors generated by classical computers into quantum states. This method calculates the inner product of these quantum states to derive the desired kernels. Embedding feature vectors into quantum states corresponds to mapping them onto a Hilbert space with dimensions raised to the power of two quantum bits (qubits). Using the kernel matrix constructed on a quantum computer, we performed a KRR, denoted as quantum KRR (QKRR).
@@
In this study, we adopted the natural parameterized quantum circuits (NPQC) Kernel, which has been demonstrated to possess performance characteristics similar to the Gaussian kernel, both theoretically and in actual hardware experiments 91, 92, 93. All computations were conducted using Scikit-qulacs 87, 94, 95. The quantum kernel was constructed in a 10-qubit space. Hyperparameters for the quantum kernel were determined through grid search. The determined parameters of NPQC kernel were = 1.5 and the repetition times of embedding 40. The regularization hyperparameter in QKRR was determined using 10 iterations of randomized search.
3 Results
@@
In Section 3.1, because we deal with many elements, we compared the dimensional efficiency of our proposed GNN-TL descriptor to well-established physics-inspired descriptors. Note that the GNN-TL descriptor can better handle complex chemical systems by exploiting the GNN-based IAP architecture.
@@
In Section 3.2, we focused on the accuracy of the GNN-TL descriptor in predicting NMR chemical shifts, which are key to understanding molecular details (, interatomic distances and bond angles). This scenario provides an ideal test for determining how well the GNN-TL descriptor works in our study.
@@
Our analysis began by comparing quantum kernel learning, in which the kernels are tested using a quantum computer emulator with traditional kernel learning methods. We then checked the accuracy of the GNN-TL descriptors across the different pretrained GNN models.
@@
Finally, we juxtaposed our GNN-TL descriptor using well-established physics-inspired descriptors. This comparison demonstrates the superiority of the proposed descriptor in terms of efficiency and accuracy. Furthermore, it highlights its potential for accurately predicting chemical properties, which is crucial for advancing research in the molecular and material sciences.
3.1 Dimensional Efficiency
| SOAP222 | FCHL19222 |
| 5 | 5,740 | 740 | 128 | 32 | 64 | 128 | 256 | 96 | 224 | ||||||||||||||
| 10 | 22,680 | 2,440 | - | - | 64 | 128 | 256 | 96 | 224 | ||||||||||||||
| 89 | 1,737,120 | 162,336 | - | - | 64 | 128 | 256 | - | - |
@@ 3.1 Dimensional Efficiency
At the atomic level, descriptors are tools designed to encode information about atoms within molecules or crystalline materials into vectors. Popular descriptors, such as SOAP and FCHL18, excel at intricately capturing the environment within an atom’s cutoff radius. Although these descriptors have achieved significant success in various accuracy benchmarks, they also present challenges due to their large dimensions. Various strategies have been developed to address these challenges,97, 98, 99, 34 including refining the descriptor itself, using principal component analysis for dimensionality reduction, and exploring NNs to encode them. In particular, Christensen et al. applied Behler’s method of the atom-centered symmetry function 100 for NN potential to discretize FCHL1833 to derive a compact and accurate FCHL19.34
@@
In Table 1, we present the scaling of the SOAP, FCHL19 and various GNN-TL descriptors in response to an increase in the number of elemental species considered. Additionally, for the QM9, QMugs,85 and MPF.2021.8 or MPtrj datasets,10 the descriptor dimensions corresponding to 5, 10, and 89 elemental species comprising each dataset are summarized, respectively. Remarkably, with an increase in the number of element types, both SOAP and FCHL19 exhibited quadratic scaling. As a snapshot, when representing five elements in the QM9 dataset, the SOAP and FCHL19 methods have dimensions of 5,740 and 740, respectively. This dimensional disparity increases with the number of elemental types. Hence, to represent the 89 elements, the dimensions increased to 1,737,120 and 162,336, respectively. These dimensions are hundreds to tens of thousands of times larger than the compact GNN-TL descriptors, which ranges from 64 to 256 dimensions. Owing to its consistent dimensionality, irrespective of the increase in elements, the GNN-TL descriptors are overwhelmingly compact.
3.2 Prediction Accuracy: NMR Chemical Shifts